OpenStudy (anonymous):

Use cramer's rule to solve the system of equations. write your answer as an ordered pair. {x-2y=-9 3x+y=1

OpenStudy (anonymous):

1 -2 3 1 that's the regular matrix switch the variable you want to solve for with the answers, so if you were to solve for x you would switch it to -9 -2 1 1

OpenStudy (anonymous):

and then you solve the matrix

OpenStudy (anonymous):

how would you solve it though, because every time I try to solve it as soon as I get it to that point I mess up?

OpenStudy (anonymous):

-9*1-1*(-2) -9-(-2) -7 so x=-7

OpenStudy (anonymous):

well if i remember correctly that's how you solve a matrix, i could be wrong though.

OpenStudy (anonymous):

So would you do the same thing for y?

OpenStudy (anonymous):

yeah for y you would do 1 -9 3 1

OpenStudy (anonymous):

so it would be 1-(-27)=28 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you should just plug the answers back into the original equation to make sure because i might be wrong

OpenStudy (anonymous):

yeah, it doesn't come out right..this is what keeps happening to me

OpenStudy (anonymous):

@sakeene831 the answer is x=-1 y=4

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

forgot to get the determinant of the equation first which is 7 after you get the determinant you replace x with -9 and 1 or y with -9 and 1 and then you divide the answer with the determinant which is 7 so in the end it's x=-7/7 and y=28/7